Please help me answer this question. This is my performance task and I will pass this tomorrow. --Suppose that five times the cube of a number equals 80 times the number. Find the number.--

Answer by Alan3354(69443)

Suppose that five times the cube of a number equals 80 times the number. Find the number
--------------
5x^3 = 80x
5x^3 - 80x = 0
5x*(x^2 - 16) = 0
x = 0
x = 4
x = -4